OpenStudy (anonymous):

Find the equation of straight line passing through point (4,5) and equally inclined to the lines 3x = 4y+7 and 5y = 12x+6

OpenStudy (anonymous):

find the angular bisector of the lines that passes through (4,5)

OpenStudy (anonymous):

7y = 9x -1

OpenStudy (anonymous):

howwwwww

OpenStudy (anonymous):

@lgbasallote @waterineyes

OpenStudy (anonymous):

clearly m1 = 3/4 and m2= 12/5 Let m be the slope of line bisceting the angle between two lines. tan A = (m-3/4) / (1+[3/4]m) = (4m - 3) /(4 + 3m) tan A = (12/5 - m) / (1+[12/5]m) = 12 -5m / 5+12m Equating we get m = 9/7 With a point (4,5) and slope m = 9/7

OpenStudy (anonymous):

Equation of line will be 7y = 9x -1

OpenStudy (anonymous):

lol i am dumb :|

OpenStudy (anonymous):

thanx